find the solution set by factoring and using the zero product rule or using the quadratic formula. 11x^2-12x=-1



Answer by htmentor(1343)

find the solution set by factoring and using the zero product rule or using the quadratic formula. 11x^2-12x=-1
=================================================
Writing in standard form we have
11x^2 - 12x + 1 = 0
This can be factored as
(11x - 1)(x - 1) = 0
This gives
x - 1 = 0 -> x = 1
11x - 1 = 0 -> x = 1/11